there are 20 people in a room singing. the mean age of 12 of the members is 18years. the mean age of the remaining 8 is 23 years. what is the mean age of the 20 members of the singing group
1
Expert's answer
2020-06-01T18:03:43-0400
"mean=\mu={\displaystyle\sum_{i=1}^nx_i\over n}"

The mean age of 12 of the members is 18 years:


"\mu_1={\displaystyle\sum_{i=1}^{12}x_i\over 12}=18""\displaystyle\sum_{i=1}^{12}x_i=12(18)=216"

The mean age of the remaining 8 is 23 years:


"\mu_2={\displaystyle\sum_{i=13}^{20}x_i\over 8}=23""\displaystyle\sum_{i=13}^{20}x_i=8(23)=184"

What is the mean age of the 20 members of the singing group?


"\mu={\displaystyle\sum_{i=1}^{20}x_i\over 20}={\displaystyle\sum_{i=1}^{12}x_i+\displaystyle\sum_{i=13}^{20}x_i\over 20}="

"={216+184\over 20}=20"

 The mean age of the 20 members of the singing group is 20 years.